"Jimmy had $5.25 in nickels and quarters. If he had 15 more nickels than quarters, how many coins of each type did he have?" HELP ME PLEASE!!!!
Answer by stanbon(75887)

Jimmy had $5.25 in nickels and quarters. If he had 15 more nickels than quarters, how many coins of each type did he have?
---------------
Let number of quarters be "x" ; Value of these is 25x cents
Number of nickels is "x+15" ; Value of these is 5(x+15)=5x+75 cents
----------
EQUATION:
value + value = 525 cents
25x + 5x+75 = 525
30x=450
x=15 (this is the number of quarters)
x+15=30 (this is the number of nickels)
-------
Cheers,
Stan H.